OpenStudy (anonymous):

The area of a rectangle is 77yd^2, and the length of the rectangle is10yd^2 less than three times the width. Find the dimensions of the rectangle.

OpenStudy (anonymous):

anybody please help!!!!

OpenStudy (anonymous):

@directrix

OpenStudy (anonymous):

@mocham

OpenStudy (anonymous):

@anybody

OpenStudy (ranga):

Assume Width = W Length = (3*W - 10yd^2) Area = L * W = W * (3*W - 10yd^2) = 77yd^2 Solve for W and then L.